Solve and check the linear equation. x+3/18=1/9+x-3/5
1
Expert's answer
2020-08-26T15:45:50-0400

In the posted question the term 'x' in the equation gets cancelled on both the sides which makes it unsolvable.


So let us consider the same equation with a small correction and solving the same as under :

(x + 3)/18 = 1/9 + (x - 3)/5


Multiplying both sides by 90 we get,


"90*\frac {(x + 3)}{18} = 90*\frac {1}{9} + 90*\frac{(x - 3)}{5}"


Hence,

5(x + 3) = 10 + 18(x - 3)

5x + 15 = 10 + 18x - 54

13x = 59


x = 59/13


So now substituting the value of x in the given equation we get,

LHS = (x + 3)/18 = (59/13 + 3)/18 = 49/117


RHS = 1/9 + (x - 3)/5 = 49/117


Since, LHS = RHS the calculated value of x = 59/13 is correct.